OpenStudy (supercm1):

Steps to get h isolated from v=1/3 x pie x r^2 x h

OpenStudy (photon336):

\[V = \frac{ 1 }{ 3 }\pi*r^{2}h\]

OpenStudy (supercm1):

Yes

OpenStudy (photon336):

to isolate this we see that we have a bunch of terms multiplied together

OpenStudy (supercm1):

To get h

OpenStudy (photon336):

now, we need to do the opposite of this which is division.

OpenStudy (photon336):

\[\frac{ V }{ \pi*r^{2} } = \frac{ 1 }{ 3 }h\]

OpenStudy (photon336):

multiply both sides by 3

OpenStudy (photon336):

Which gives us \[\frac{ 3V }{ \pi*r^{2} } = h \]

OpenStudy (supercm1):

Thanks